# PRISM: Prompt-Refined In-Context System Modelling for Financial Retrieval

### v1

@@ Abstract
With the rapid progress of large language models (LLMs), financial information retrieval has become a critical industrial application. Extracting task-relevant information from lengthy financial filings is essential for both operational and analytical decision-making. The FinAgentBench dataset finagentbench formalizes this problem through two tasks: document ranking and chunk ranking. We present PRISM, a training-free framework that integrates refined system prompting, in-context learning (ICL), and a lightweight multi-agent system. Each component is examined extensively to reveal their synergies: prompt engineering provides precise task instructions, ICL supplies semantically relevant few-shot examples, and the multi-agent system models coordinated scoring behaviour. Our best configuration achieves an NDCG@5 of 0.71818 on the restricted validation split. We further demonstrate that PRISM is feasible and robust for production-scale financial retrieval. Its modular, inference-only design makes it practical for real-world use cases. The source code is released at https://bit.ly/prism-ailens.
PRISM: Prompt-Refined In-Context System Modelling
for Financial Retrieval
[…]
@@ 1 Introduction
Large language model (LLM) based Information Retrieval (IR) has emerged as a transformative technology in recent years, particularly within the financial domain zhu2023large. Lengthy and complex financial documents such as annual reports and cash flow statements often require substantial human effort to extract key information, and the process remains highly prone to error zhao2024optimizing. As a result, research focus has shifted toward evidence-based information retrieval, where systems are expected to provide verifiable sources of answers through a Retrieval-Augmented Generation (RAG) workflow choi2025finder. Despite these advances, significant challenges persist due to the dense and domain-specific nature of financial texts. LLMs often struggle to accurately identify both the “what” and the “where” of correct answers when responding to user queries. Recently, FinAgentBench was introduced as a large scale benchmark dataset to study and evaluate two critical components: document ranking, which determines the relevance of financial documents, and chunk ranking, which identifies the most relevant segments within a document finagentbench. This two-step approach aims to produce accurate and evidence-grounded responses. Building on recent advances in LLMs, we investigate how OpenAI’s Generative Pre-trained Transformer (GPT), one of the most advanced commercially available LLMs islam2023financebench, can help address these challenges. We propose Prompt-Refined In-Context System Modelling (PRISM), a training-free framework that integrates precise system prompt engineering, in-context learning (ICL) augmentation, and multi-agent system modelling.
Figure 1: Overview of the proposed PRISM framework.
We begin by conducting exploratory data analysis (EDA) at both the document and chunk levels to uncover structural and semantic patterns within the data. The insights obtained from this analysis guide the design of prompt templates, ICL augmentation, and agent routing logic. Systematic prompt engineering is then applied to construct reasoning-oriented prompts that enhance the model’s ability to rank documents and text segments effectively. Next, we introduce ICL augmentation, which leverages OpenAI’s latest text embedding models to store question–answer pairs in a vector database and retrieve them as few-shot examples during inference on the test set. Finally, these modules are synchronized within a multi-agent framework, where specialized agents collaborate to improve ranking performance and overall system reliability. The contributions of this paper are threefold:
@@
    We propose PRISM, a framework for document and text chunk ranking that requires no additional training or fine-tuning of LLMs.
@@
    We conduct extensive experiments and ablation studies to evaluate the effects and interactions of the different modules within PRISM.
@@
    We perform reproducibility and feasibility studies to assess the consistency and deployability of PRISM in real-world financial IR.
2 Related Works
@@
Prompt Engineering. Recent studies have shown that retrieval performance is highly sensitive to how an LLM is instructed zhang2025scientific. The initial breakthrough in few-shot learning demonstrated that LLMs can perform tasks using in-context exemplars without gradient updates fewshotlearners. However, traditional few-shot approaches struggled with complex reasoning tasks. This limitation was addressed by Chain-of-Thought (CoT) prompting, which enhances reasoning by guiding LLMs through intermediate natural language reasoning steps chainofthought. The Reasoning and Acting (ReAct) framework further improves reliability by allowing LLMs to plan and execute reasoning steps iteratively, reducing hallucination reactLLM. Building on this, the Tree-of-Thoughts (ToT) framework enables LLMs to explore and evaluate multiple reasoning paths to select the next step treeofthought.
@@
LLMs for Ranking. LLMs have also demonstrated strong potential as zero-shot rankers in IR, typically categorized into pointwise, pairwise, and listwise approaches LLMZeroShotRankers. Listwise methods, such as RankGPT, are often preferred for their balance between effectiveness and efficiency, but they face key challenges including limited input length and sensitivity to document order chatgptLLMRerank. TourRank introduces a multi-stage grouping and tournament strategy to handle large candidate sets and improve robustness to input ordering tourrankLLM. Similarly, Pairwise Ranking Prompting simplifies the comparison process, allowing moderately sized LLMs to achieve state-of-the-art performance LLMTextRerankPairwiseRanking. Generally, well-instructed LLMs have demonstrated superior ranking performance compared to state-of-the-art supervised systems on major benchmarks such as TREC-DL and BEIR chatgptLLMRerank. However, existing unsupervised methods remain constrained by LLMs’ token limits when applied to long and dense financial texts zhao2024docmath.
@@
In-Context Learning (ICL). In-context learning (ICL) enables LLMs to adapt to new tasks using input–output examples in the prompt, without gradient updates. Recent work views this mechanism as implicit Bayesian inference, where the model infers a latent concept that explains the observed examples ICLBayesianInference. Empirically, ICL accuracy scales with both the number and length of examples but is not strictly tied to correct input–label mappings. Randomly replacing labels only marginally affects performance WhatMakesICLWork, suggesting that ICL primarily benefits from contextual cues defining the label space, input distribution, and output format. The emergence of ICL has also been linked to properties of pre-training data such as temporal burstiness and diverse, low-frequency classes ICLDataDistributionalProperties. Meta-in-context learning aims to refine the model’s implicit priors and adaptive strategies presenting multiple tasks sequentially metaICL. However, there are limited research on embedding-based ICL retrieval for financial IR with commercial LLMs.
@@
Agentic Information Retrieval. Agentic information retrieval represents a new paradigm where LLM agents dynamically manage information access through iterative cycles of observation, reasoning, and action, distinguishing it from traditional IR architectures zhang2025agenticinformationretrieval. It is commonly implemented using multi-agent system (MAS), which coordinate specialized agents for reflection, planning, and tool use to achieve collective intelligence and outperform single-agent approaches singh2025agenticretrievalaugmentedgenerationsurvey. LLM agents have also been applied to ranking tasks, demonstrating that retrieval and ranking agents can effectively coexist xu-etal-2025-iagent. Despite its potential, Agentic IR faces challenges in coordinating complex MAS interactions, reducing the high inference cost of LLMs, and evaluating dynamic agentic behaviours in financial retrieval contexts.
3 FinAgentBench Dataset
@@
FinAgentBench is a large-scale dataset for financial information retrieval (IR), comprising 2023–2024 financial documents from the EDGAR database finagentbench. It supports two tasks: given a user query, (i) ranking five document types (10-K, 10-Q, 8-K, DEF 14A, and Earnings Transcripts) that are most likely to contain the answer, and (ii) ranking text chunks within a document based on their relevance. The dataset includes a training set of 4,986 document-ranking samples and 18,855 chunk-ranking samples, and a validation set of 200 samples for each task.
Document Ranking Data Analysis. We first conduct a systematic evaluation across all document types in FinAgentBench. Table 1 summarizes the distribution of document types across ranking positions. Documents such as 10-K and Earnings Transcripts appear more frequently in higher ranks, while 8-K filings are more evenly distributed across lower ranks. This suggests that different document types vary in information density and relevance, which may influence retrieval difficulty in downstream tasks. Table 2 shows the six most frequent keywords found in Rank 1 documents for each type, after removing stopwords. DEF 14A filings are dominated by high-ratio keywords such as dependency and concentration, while 8-K filings focus more on compensation-related terms. 10-Q filings contain lower keyword ratios, reflecting greater variation in language use. In contrast, Earnings Transcripts and DEF 14A documents show stronger keyword concentration, suggesting that certain document types have more distinctive language patterns that can help guide retrieval. These findings highlight how differences in vocabulary and document relevance can affect ranking performance and point to the need for adaptive retrieval strategies.
@@
Doc. Type Rank 1 Rank 2 Rank 3 Rank 4 Rank 5 DEF 14A 573 749 1013 1212 1439 10-K 2402 1546 507 207 324 10-Q 554 1558 1427 866 581 8-K 97 407 1091 1808 1583 Earnings 1360 726 948 893 1059
Table 2: Top 6 Rank 1 keywords by document type.
Doc. Type Keyword Ratio Keyword Ratio DEF 14A dependency 88.40% concentration 88.24% exist 87.73% risks 80.51% market 60.77% share 57.12% 10-K evolved 40.98% recurring 40.57% time 39.22% ratio 39.40% reporting 38.19% period 36.01% 10-Q quarter 20.00% recurring 6.29% evolved 6.12% time 5.99% ratio 5.97% revenue 4.02% 8-K compensation 95.45% award 41.74% burn 41.74% manage 41.54% availability 40.60% share 27.21% Earnings asked 87.68% questions 86.56% metrics 78.26% customer 74.62% guidance 52.97% offered 52.80%
Chunk Ranking Data Analysis. We conducted a frequency analysis to better understand chunk-level characteristics in the dataset. As shown in Figure 2, relevant chunks are generally longer and more information dense than irrelevant ones. The wider interquartile ranges and higher maximum values, along with the long-tail distribution, indicate substantial variability in chunk length. This suggests that a dynamic retrieval and ranking method is needed to handle both typical and unusually long chunks to avoid truncation and processing inefficiencies. Moreover, the relatively small proportion of relevant chunks highlights the fine-grained and selective nature of financial retrieval. Hence, these findings indicate that models should focus on reasoning with information-rich chunks while using adaptive strategies to manage large variations in chunk length and maintain ranking accuracy.
@@ 4 PRISM
Our EDA shows that relevant answers are concentrated in specific document types and in a small minority of information-dense chunks. These findings guided the design of PRISM, which leverages strong domain priors, dynamic ICL retrieval and multi-agent system to improve ranking relevance.
@@ 4.1 Prompt Engineering
This foundational module evaluates how logical constraints and domain-specific priors embedded in prompts influence LLM reasoning quality. We introduced logical scaffolding inspired by ReAct, CoT, and ToT to promote explicit reasoning before producing final answers. Four prompt variants, $P_{1}$$P_{4}$, were systematically designed to assess the impact of incremental modifications and progressively enhance reasoning structure.
The $P_{1}$ variant followed a ReAct-style design with corpus-informed guidance to capture structural regularities in document and chunk distributions. $P_{2}$ added quantitative domain priors, such as keyword ratios, to complement conceptual cues but showed minimal improvement over $P_{1}$. $P_{3}$ incorporated explicit reasoning scaffolds from CoT and ToT to better handle long, information-dense chunks, while $P_{4}$ streamlined the prompt to be more straightforward by reducing hallucination risk, and enforcing stricter output formats for downstream re-ranking. Structural refinements in $P_{3}$$P_{4}$ outperformed content-based augmentations, with $P_{4}$ achieving the best overall results. Prompt templates are provided in Section A.8 of the Appendix.
@@ 4.2 In-Context Learning Sample Retrieval
The second module enhances the system with few-shot learning through a simplified RAG pipeline. Our EDA shows that lexical cues are unevenly distributed across SEC document types, with certain types dominating the top ranks and displaying distinct keyword patterns. Incorporating examples that reflect these domain priors helps LLMs identify relevant documents earlier and reason more consistently. At the chunk level, relevant spans are generally longer and more information-dense, increasing the risk of over-elaboration and error propagation. Providing a few well-chosen examples of relevant cases gives the model a stable reference point, reducing hallucinations and improving ranking stability in long-context settings.
We embedded all query–document/chunk pairs and their ground-truth rankings using OpenAI’s text-embedding-3 model and built a FAISS vector store. During inference, each query retrieves its top-$k$ most semantically similar exemplars, which are formatted into coherent few-shot examples and prepended to the system prompt. This makes the final prompt dynamic, adapting to the semantic likelihood of each query rather than relying on a static template. Consistent with prior work, we find that ICL performs best when added samples are semantically close to the target query.
4.3 Multi-Agent System (MAS) Modelling
The third module introduces a structured MAS to decompose reasoning into specialized agent roles. This design aims to reduce cognitive load, improve calibration, and mitigate hallucinations in the ranking tasks. By enforcing role specialization and graph-level coordination, the MAS constrains and regularizes LLM behaviour for more consistent outcomes. We adopt a state-graph framework with architectural variations focused primarily on chunk ranking, a task that is sensitive to error propagation and reasoning collapse. Document ranking agents consist of a Question Analyzer and five domain-specific Document Experts that are remain fixed across all configurations to ensure consistent document ranking and isolate chunk-level effects.
As for chunk ranking, the MAS defines four progressively simplified architectures, $A_{1}$$A_{4}$. $A_{1}$ uses a direct scoring ensemble of four parallel agents that independently evaluate candidate chunks before reaching a consensus. $A_{2}$ extends this into a three-stage pipeline with a noise filter, candidate selector, and scoring ensemble, but tends to suffer from over-filtering and cascading errors. $A_{3}$ simplifies the structure to two stages which has a quick filter followed by three scoring agents to balance depth and stability. Finally, $A_{4}$ adopts a minimal configuration with two scoring agents, reducing interaction complexity and error propagation while maintaining consistent performance.
Table 3: Ablation studies of document and chunk ranking configs. across agentic, non-agentic and hybrid workflows.

### v2

# PRISM: Prompt-Refined In-Context System Modeling for Financial Retrieval

@@ Abstract
With the rapid progress of large language models (LLMs), financial information retrieval has become a critical industrial application. Extracting task-relevant information from lengthy financial filings is essential for both operational and analytical decision-making. We present PRISM, a training-free framework that integrates refined system prompting, in-context learning (ICL), and lightweight multi-agent coordination for document and chunk ranking tasks. Our primary contribution is a systematic empirical study of when each component provides value: prompt engineering delivers consistent performance with minimal overhead, ICL enhances reasoning for complex queries when applied selectively, and multi-agent systems show potential primarily with larger models and careful architectural design. Extensive ablation studies across FinAgentBench, FiQA-2018, and FinanceBench reveal that simpler configurations often outperform complex multi-agent pipelines, providing practical guidance for practitioners. Our best configuration achieves an NDCG@5 of 0.71818 on FinAgentBench, ranking third while being the only training-free approach in the top three. We provide comprehensive feasibility analyses covering latency, token usage, and cost trade-offs to support deployment decisions. The source code is released at https://bit.ly/prism-ailens.
@@ 1 Introduction
Figure 1: Overview of the PRISM framework. User queries are processed through three independent components: (1) prompt engineering encodes domain priors and reasoning scaffolds, (2) ICL retrieval augments prompts with semantically similar examples from the training set, and (3) optional multi-agent coordination decomposes ranking into specialized agent roles. The system outputs ranked documents and chunks for financial information retrieval.
Large language model (LLM) based Information Retrieval (IR) has emerged as a transformative technology in the financial domain (Zhu et al., 2023). Extracting key information from lengthy financial documents remains labor-intensive and error-prone (Zhao et al., 2024b), driving research toward evidence-based retrieval where systems provide verifiable answers through Retrieval-Augmented Generation (RAG) workflows (Choi et al., 2025a). Despite these advances, challenges persist due to the dense and domain-specific nature of financial texts. FinAgentBench was recently introduced to evaluate two tasks: document ranking to determine the relevance of financial documents, and chunk ranking to identify the most relevant segments within a document (Choi et al., 2025b). Building on recent advances, we investigate how OpenAI’s Generative Pre-trained Transformer (GPT) (Islam et al., 2023) can address these challenges. We propose Prompt-Refined In-Context System Modeling (PRISM), a training-free framework that integrates precise system prompt engineering, in-context learning (ICL) augmentation, and multi-agent system modeling. Our contributions are:
@@
    We present PRISM, a training-free framework that integrates prompt engineering, in-context learning, and multi-agent coordination for financial document and chunk ranking, evaluated across FinAgentBench, FiQA-2018 and FinanceBench datasets.
@@
    We conduct systematic ablation studies that examine how prompt design, ICL, and multi-agent coordination interact, identifying the conditions under which each component provides value and showing that simpler non-agentic configurations frequently outperform more complex multi-agent pipelines, particularly for fine-grained chunk ranking tasks.
@@
    We provide the first latency, token usage, and cost analysis, establishing reproducible baselines for evaluating accuracy-efficiency trade-offs in training-free financial IR systems.
2 Related Work
@@
Prompt Engineering. Recent studies have shown that retrieval performance is highly sensitive to how an LLM is instructed (Zhang et al., 2025b). The initial breakthrough in few-shot learning demonstrated that LLMs can perform tasks using in-context exemplars without gradient updates (Brown et al., 2020). However, traditional few-shot approaches struggled with complex reasoning tasks. This limitation was addressed by Chain-of-Thought (CoT) prompting, which enhances reasoning by guiding LLMs through intermediate natural language reasoning steps (Wei et al., 2022). The Reasoning and Acting (ReAct) framework further improves reliability by allowing LLMs to plan and execute reasoning steps iteratively, reducing hallucination (Yao et al., 2023b). Building on this, the Tree-of-Thoughts (ToT) framework enables LLMs to explore and evaluate multiple reasoning paths to select the next step (Yao et al., 2023a).
@@
LLMs for Ranking. LLMs have also demonstrated strong potential as zero-shot rankers in IR, typically categorized into pointwise, pairwise, and listwise approaches (Hou et al., 2024). Listwise methods, such as RankGPT, are often preferred for their balance between effectiveness and efficiency, but they face key challenges including limited input length and sensitivity to document order (Sun et al., 2024). TourRank introduces a multi-stage grouping and tournament strategy to handle large candidate sets and improve robustness to input ordering (Chen et al., 2025). Similarly, Pairwise Ranking Prompting introduces pairwise comparison of output with linear complexity (Qin et al., 2024). Generally, well-instructed LLMs have demonstrated superior ranking performance compared to state-of-the-art supervised systems on major benchmarks such as TREC-DL and BEIR (Sun et al., 2024). However, existing unsupervised methods remain constrained by LLMs’ token limits when applied to long and dense financial texts (Zhao et al., 2024a).
@@
In-Context Learning (ICL). In-context learning (ICL) enables LLMs to adapt to new tasks using input-output examples in the prompt, without gradient updates. Recent work views this mechanism as implicit Bayesian inference, where the model infers a latent concept that explains the observed examples (Xie et al., 2022). Empirically, ICL accuracy scales with both the number and length of examples but is not strictly tied to correct input-label mappings. Randomly replacing labels only marginally affects performance (Min et al., 2022), suggesting that ICL primarily benefits from contextual cues that define the label space, input distribution, and output format. The emergence of ICL has also been linked to properties of pre-training data such as temporal burstiness and diverse, low-frequency classes (Chan et al., 2022). Meta-ICL aims to refine the model’s implicit priors and adaptive strategies by presenting multiple tasks sequentially (Coda-Forno et al., 2023). However, embedding-based ICL retrieval for financial IR with commercial LLMs remains underexplored.
@@
Agentic Information Retrieval. Agentic information retrieval represents a new paradigm where LLM agents dynamically manage information access through iterative cycles of observation, reasoning, and action, distinguishing it from traditional IR architectures (Zhang et al., 2025a). It is commonly implemented using multi-agent systems (MAS), which coordinate specialized agents for reflection, planning, and tool use to achieve collective intelligence and outperform single-agent approaches (Singh et al., 2025). LLM agents have also been applied to ranking tasks, demonstrating that retrieval and ranking agents can effectively coexist (Xu et al., 2025). Despite its potential, Agentic IR faces challenges in coordinating complex MAS interactions, reducing the high inference cost of LLMs, and evaluating dynamic agentic behaviors in financial retrieval contexts.
3 Datasets
We evaluate PRISM on three financial datasets that cover complementary task characteristics: ranking granularities (document vs. chunk vs. passage), task types (ranking vs. question answering), and reasoning demands (lexical matching vs. numerical analysis).
FiQA-2018 (Maia et al., 2018) is a financial opinion QA dataset from the BEIR benchmark (Thakur et al., 2021), containing questions sourced from financial microblogs, news headlines, and reports. The task requires passage reranking to identify relevant answers. NDCG@10 and Recall@100 are used as evaluation metrics, comparing against baselines such as BM25 and cross-encoder rerankers.
FinanceBench (Islam et al., 2023) is a question-answering benchmark requiring numerical reasoning over financial documents. Unlike ranking tasks, it tests whether systems can extract correct factual answers from SEC filings. We evaluate under oracle context conditions (relevant passages provided) to isolate the reasoning component from retrieval.
@@ 3 Datasets
FinAgentBench (Choi et al., 2025b) is a large-scale dataset for financial IR comprising 2023–2024 SEC filings from the EDGAR database. It supports two tasks: (i) ranking five document types (10-K, 10-Q, 8-K, DEF 14A, Earnings Transcripts) by relevance to a query, and (ii) ranking text chunks within documents. The dataset includes 4,986 document-ranking and 18,855 chunk-ranking training samples, with 200 validation samples per task.
3.1 Exploratory data analysis (EDA) on FinAgentBench
@@
Doc. Type Rank 1 Rank 2 Rank 3 Rank 4 Rank 5 Doc. Type Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
DEF 14A 573 749 1013 1212 1439 10-Q 554 1558 1427 866 581
10-K 2402 1546 507 207 324 8-K 97 407 1091 1808 1583
Earnings 1360 726 948 893 1059
Table 2: Top 6 keywords appearing in queries where the respective document type was retrieved at Rank 1. The ratio indicates the frequency of the keyword relative to the entire corpus.
DEF 14A 10-K 10-Q 8-K Earnings
Keyword Ratio Keyword Ratio Keyword Ratio Keyword Ratio Keyword Ratio
dependency 88.40% evolved 40.98% quarter 20.00% compensation 95.45% asked 87.68%
concentration 88.24% recurring 40.57% recurring 6.29% burn 41.74% questions 86.56%
exist 87.73% ratio 39.40% evolved 6.12% award 41.74% metrics 78.26%
risks 80.51% time 39.22% time 5.99% manage 41.54% customer 74.62%
market 60.77% reporting 38.19% ratio 5.97% availability 40.60% guidance 52.97%
share 57.12% period 36.01% revenue 4.02% share 27.21% offered 52.80%
Document Ranking Data Analysis. We first conduct a systematic evaluation across all document types in FinAgentBench. Table 1 summarizes the distribution of document types across ranking positions. Documents such as 10-K and 8-K appear more frequently in higher and lower ranks respectively, suggesting that document types vary in information density and relevance. Table 2 shows the six most frequent keywords found in Rank 1 documents for each type. These differences in vocabulary concentration across document types highlight the need for adaptive retrieval strategies; a detailed keyword analysis is provided in Appendix A.1.1.
Chunk Ranking Data Analysis. We conducted a frequency analysis to better understand chunk-level characteristics in the dataset. As shown in Figure 2, relevant chunks are generally more information dense than irrelevant ones. The wider interquartile range (IQR) and higher maximum values, along with the long-tail distribution, suggest that adaptive retrieval methods are needed to handle substantial variability in chunk length. Moreover, the relatively small proportion of relevant chunks highlights the fine-grained and selective nature of financial retrieval. Overall, our EDA shows that relevant answers are concentrated in specific document types and in a small minority of information-dense chunks, requiring models to reason and manage large variations in chunk length.
FiQA-2018 and FinanceBench. We performed EDA on FiQA-2018 and FinanceBench to examine relevance sparsity, query-document length characteristics, and lexical regularities that influence retrieval difficulty. Comprehensive analyses are deferred to Appendix A.2 and Appendix A.3.
@@ 4 PRISM
Guided by the EDA findings above, PRISM integrates three components whose interactions are illustrated in Figure 1: (1) prompt engineering that encodes domain priors and reasoning scaffolds, (2) ICL retrieval that dynamically augments prompts with semantically relevant examples, and (3) multi-agent coordination that decomposes ranking into specialized agent roles.
@@ 4.1 Prompt Engineering
Prompt engineering in PRISM is an integrated system module that governs how LLMs process ranking tasks. This module encodes domain-specific priors derived from our EDA, such as document-type distributions and keyword patterns, into structured instructions that guide the model’s reasoning. We introduce logical scaffolding inspired by ReAct, CoT, and ToT to promote explicit reasoning before producing final answers. Four prompt variants, $P_{1}$$P_{4}$, were systematically designed to assess the impact of incremental modifications and progressively enhance reasoning structure. The $P_{1}$ variant followed a ReAct-style design with corpus-informed guidance to capture structural regularities in document and chunk distributions. $P_{2}$ added quantitative domain priors, such as keyword ratios, to complement conceptual cues but showed minimal improvement over $P_{1}$. $P_{3}$ incorporated explicit reasoning scaffolds from CoT and ToT to better handle long, information-dense chunks, while $P_{4}$ streamlined the prompt to be more straightforward by reducing hallucination risk and enforcing stricter output formats for downstream re-ranking. Structural refinements in $P_{3}$$P_{4}$ outperformed content-based augmentations, with $P_{4}$ achieving the best overall results, confirming that reasoning clarity outweighs informational density. Prompt templates are provided in Appendix A.4.
@@ 4.2 In-Context Learning Sample Retrieval
The second module enhances the system with few-shot learning through a simplified RAG pipeline. As our EDA reveals uneven lexical distributions and varying information density across document types and chunks (Tables 12), incorporating well-chosen examples that reflect these domain priors gives LLMs a stable reference point, helping them identify relevant documents earlier, reason more consistently, and reduce hallucinations in long-context settings.
Exemplar Store Construction. We first construct an offline exemplar store from the training dataset. Each training sample consists of a query paired with documents or chunks and their ground-truth relevance rankings. We embed all query-document/chunk pairs using OpenAI’s text-embedding-3 model and index them in a FAISS vector store. This store remains fixed and serves as the source for retrieving few-shot examples.
Inference-Time Retrieval. During inference, each new query is embedded using the same model and the top-$k$ most semantically similar exemplars are retrieved from the FAISS store. These exemplars, including example queries and their ground-truth rankings, are formatted as few-shot demonstrations and prepended to the system prompt, making each prompt dynamically adapted to the query. Further details on store separation and retrieval dynamics are provided in Appendix A.6.
4.3 Multi-Agent System (MAS) Modeling
The third module introduces a structured MAS to decompose reasoning into specialized agent roles, aiming to reduce cognitive load, improve calibration, and mitigate hallucinations through role specialization and graph-level coordination. Despite mixed empirical results, we include MAS to systematically study the conditions under which multi-agent coordination provides value (larger models, document-level tasks) versus where it falls short (smaller models, chunk-level tasks with error propagation). This nuanced understanding is a key contribution of our work.
Architectural Variants. For chunk ranking, the MAS defines four progressively simplified architectures ($A_{1}$$A_{4}$): $A_{1}$: Direct scoring ensemble with four parallel agents (CEO, Financial Analyst, Operations Manager, Risk Analyst) that independently evaluate candidate chunks before reaching consensus through score averaging. $A_{2}$: Three-stage pipeline with Noise Remover $\rightarrow$ Candidate Selector $\rightarrow$ scoring ensemble (Relevance Scorer, Contextual Reasoner, Evidence Extractor, Diversity Agent). This deeper architecture tends to suffer from over-filtering and cascading errors. $A_{3}$: Two-stage pipeline with Quick Filter $\rightarrow$ three scoring agents (Relevance Scorer, Contextual Reasoner, Evidence Extractor), balancing depth with stability. $A_{4}$: Minimal configuration with two scoring agents (Financial Analyst, Risk Analyst), reducing interaction complexity and error propagation. Detailed explanations of each variant and its workflow are included in Appendix A.7.
Table 3: Ablation studies of document and chunk ranking configurations across different workflows.